OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5). 102 52 22 −1.5

hartnn (hartnn):

just put x=5 in f(x)

OpenStudy (anonymous):

wouldnt dat be -1.5

hartnn (hartnn):

nopes, what is \(\huge 2\times 5^2+2=?\)

OpenStudy (anonymous):

i seen this question twice today lol f(5) = 2(5)^2 + 2

hartnn (hartnn):

5^2 = ?

OpenStudy (anonymous):

52

OpenStudy (anonymous):

thats the answer for the whole thing 52 is correct

hartnn (hartnn):

yes, 52 is correct :)

OpenStudy (anonymous):

thanx for helping me